The perimeter of a rectangle is 54. If the length is three more than twice the width, what are the dimensions of the rectangle?
Answer by checkley77(12844)

L=2W+3
2L+2W=54
2(2W+3)+2W=54
4W+6+2W=54
6W=54-6
6W=48
W=48/6
W=8 ANS.
L=2*8+3
L=16+3=19 ANS.
PROOF:
2*19+2*8=54
38+16=54
54=54